OpenStudy (anonymous):

20 girls left, now there are 2 boys for each girl. Then 20 of the boys left too. There were now 3 times as many girls as boys. how many boys and girls were at the beginning of theparty?

OpenStudy (callisto):

Let x and y be the number of boys and girls at the beginning of the party. From ''20 girls left, now there are 2 boys for each girl.'', that means the number of boy is twice of that of girls. Number of girls left = y - 20 (since 20 of them left) number of boy is twice of that of girls => number of boys = 2 x number of girls => x = 2 (y-20) ----(1) From ''Then 20 of the boys left too. There were now 3 times as many girls as boys. '' Number of boys left = x-20 There were now 3 times as many girls as boys. => Number of girls = 3 x number of boys => (y-20) = 3 (x-20) ---(2) Now, solve equation (1) and (2).

OpenStudy (anonymous):

there were 40 boys and 33 girls at he beginning of the part???? still dont understand!!!

OpenStudy (anonymous):

can someone help?

OpenStudy (callisto):

That's not the answer I've got... Which part you don't understand?

OpenStudy (anonymous):

for the first equation i get 20=y Does it means that there were 20 girls left after 20 were gone? therefore there were 40 girls at the beginning of the party/

OpenStudy (anonymous):

is it right?

OpenStudy (callisto):

Nope, it doesn't seem right! How do you get 20=y from the first equation?

OpenStudy (anonymous):

x=2(y-20) x=2y-40 40x=2y all divided by 2 20x=y

OpenStudy (callisto):

For the step :40x=2y all divided by 2, it is not right. x=2y-40 You can only add 40 to both sides, then divide both sides by two to make y as the subject. But you don't have to do so actually.

OpenStudy (callisto):

It is a equation in two unknowns, so, basically, you cannot solve it alone.

OpenStudy (anonymous):

so for equatin one this is all th eunswer: x=2y-40?

OpenStudy (anonymous):

for equatin 2 i have: y-20=3x-60

OpenStudy (callisto):

Yup! Now, you need to solve these two equations. Do you know how to solve it?

OpenStudy (anonymous):

no

OpenStudy (callisto):

Have you learnt how to solve simultaneous questions yet?

OpenStudy (anonymous):

if i did it was looong time ago. this should be easy my class is called: foundations of math x teaching. so, i guess anyone should be able to solve this. but no me...( its hard x me

OpenStudy (callisto):

Hmm.. For this question. You get two question so far. First one x= 2y-40, the second one, y-20 = 3x-60. You can use the method of substitution to solve it. That is substituting equation one into two. I meant replace x by 2y-40 in the second equation so that the equation would be in one variable and you can solve y.

OpenStudy (anonymous):

The ratio of boys to girls after 20 girls left is:\[\frac{B}{G-20}=2\]The ratio of girls to boys after an additional 20 girls left is:\[\frac{G-20}{B-20}\text{= }3\]Solve the above equations for B and G.\[\{B=24,G=32\} \]

OpenStudy (anonymous):

is it: y=6y-180+20 y=6y-60 divided by 6 y=negative 10.....is no right..i know. what is it?

OpenStudy (callisto):

The first step is wrong. Remember the second equation is y - 20 = 3x - 60 Sub. x= 2y-40 into the above equation: y - 20 = 3 ( 2y-40 ) - 60 <- this is the right one.

OpenStudy (anonymous):

My bad. "The ratio of girls to boys after an additional 20 girls left is:" should have been written: "The ratio of girls to boys after an additional 20 boys left is:"

OpenStudy (anonymous):

to Callisto: yes i did that but then i did this: y-20=3(2y-40)-60 y-20=6y-120-60

OpenStudy (anonymous):

to callisto; IS IT; y-20=(6y-120)-60

OpenStudy (anonymous):

32 girls take away 20 girls is 12 girls compared to 24 boys; there are now twice as many boys as girls. 24 boys take away 20 boys leaves 4 boys compared to 12 girls; there are now three times as many girls as boys.

OpenStudy (callisto):

Yup, you're on the right track!

OpenStudy (callisto):

So, do the additions/subtractions for like terms. Isolate y also to solve y.

OpenStudy (anonymous):

y-20=(6y-120)-60 plese help on this whats the nxt step..

OpenStudy (callisto):

y-20=(6y-120)-60 y-20=6y-120-60 Add 20 to both sides y-20 + 20 =6y-120-60 + 20 What do you get?

OpenStudy (anonymous):

y=6y-160

OpenStudy (callisto):

Then, subtract both sides by 6y, what do you get?

OpenStudy (anonymous):

y=154

OpenStudy (callisto):

Why?

OpenStudy (anonymous):

y=6y-180+20...now subtract both sides by 6 y-6=-160 y=154

OpenStudy (callisto):

Subtract both sides by 6 doesn't help you cancel 6y. You need to subtract 6y from both sides.

OpenStudy (anonymous):

y=-160? please ......tell me more i need to resolve it x my class is in 20 minutes.

OpenStudy (callisto):

y=6y-180+20 subtract both sides by 6y y - 6y=6y-180+20 - 6y What do you get after simplifying it?

OpenStudy (anonymous):

y=32

OpenStudy (anonymous):

5y=-160 then i divided all by 5 then i got y=-32 but i know it can b a neg. number. ist right -32?

OpenStudy (callisto):

-32 is not right. You made a mistake in your first step. y-6y = -5y, not 5y

OpenStudy (callisto):

That's the reason why you get a negative value for y.

OpenStudy (anonymous):

i got it now! thank u very very much x all ur patience with me . i would try to resolve the rest!

OpenStudy (callisto):

You're welcome :) Tell me if you have any problems for this question, I would try to help if I can :)

OpenStudy (anonymous):

answer: 32 girls and 24 boys at beginning of party. is it right?

OpenStudy (anonymous):

Yes. Refer to my post, 24 posts back.